OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (ghazi):

to find the slope either try to write equation in X/A+Y/B=1 form or else try to write y=mx +c...you'll get slope as well as y intercept

OpenStudy (ghazi):

y= 5/3x+2....